DTO, Inc., has sales of $40 million, total assets of $24 million, and total debt of $5 million. a. If the profit margin is 8 percent, what is the net income? b. What is the ROA? c. What is the ROE?
Answer #1

A B C D E Sales Total assets Total debt Profit margin a) net income 3.2 b) ROA 13.33% C) ROE 16.84%

A B C D Sales Total assets Total debt Profit margin 0.08 a) net income =D5*D8 b) ROA =D10/D6 C) ROE =D10/(D6-07)